If I am driving a 1000 kg car at a speed of 15 m/s at the bottom of a hill and let the car coast (NO
more gas given to the car) how high up the hill will the car be able to reach? SHOW YOUR WORK
1
Expert's answer
2020-11-19T09:20:01-0500

If we assume zero friction, then all kinetic energy of car at the bottom of the hill converts to the potential energy of its at the top of the hill:


"\dfrac{mv^2}{2}=mgh"

where "m = 1000kg" is the mass of the car, "g = 9.81m\/s^2" is the gravitational acceleration, "h" is the height of the hill, and "v = 15m\/s" is the speed of the car at the bottom of the hill. Thus, obtain:


"h = \dfrac{v^2}{g}\\
h=\dfrac{15^2}{9.81} \approx 22.9m"

Answer. 22.9 m.